Circular No. 6884

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980425 P. Soffitta, M. Feroci, and L. Piro, Istituto Astrofisica Spaziale, CNR,

Rome; J. in ’t Zand and J. Heise, Space Research Organization of the Netherlands (SRON), Utrecht; L. Di Ciolo, BeppoSAX Science Operation Center, Rome; J. M. Muller, BeppoSAX Science Data Center, Rome, and SRON; and E. Palazzi and F. Frontera, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna, write: “The BeppoSAX Gamma-Ray Burst Monitor was triggered by a γ-ray burst on Apr. 25.90915 UT. The γ-ray burst appears to be single-peaked and not structured. A preliminary analysis reveals a duration of ∼ 30 s and a peak intensity of 350 counts/s in the energy band 40–700 keV. The burst was also detected in Wide Field Camera no. 2 with comparable duration and a peak intensity of ∼ 3 Crab (2–28 keV). The position derived by the WFC image is α = 19^{h}34^{m}54^{s}, δ = −52^{o}49′.9 (equinox 2000.0), with an error radius of 8′ that includes uncertainties due to a non-optimum attitude control configuration. A follow-up observation with BeppoSAX narrow-field instruments is in progress.”

COMET 1998 H1

Additional precise positions:

1998 UT α_{2000} δ_{2000} m_{1} Observer
Apr. 26.69787 15^{h}27^{m}37.^{s}49 +11^{o}55′45′′.8 12.4 Sugie
26.70065 15 27 36.37 +11 56 11.8  _{”}
26.70620 15 27 34.18 +11 57 02.9  _{”}
26.79369 15 27 00.29 +12 10 40.2  Tichá
26.79503 15 26 59.75 +12 10 52.1  _{”}
26.79635 15 26 59.25 +12 11 05.0  _{”}
26.79700 15 26 59.03 +12 11 10.9  _{”}
26.79990 15 26 57.86 +12 11 39.0 12.5  _{”}

A. Sugie (Dynic Astronomical Obs.). 0.60-m f/4.0 reflector + CCD. Tail

at least 10′ long toward south-southwest. J. Tichá and M. Tichý (Kleť). 0.57-m reflector + CCD. Measurer Tichý.

SUPERNOVA 1998aq IN NGC 3982 Precise position by G. Masi, Ceccano, Italy, for SN 1998aq (cf. IAUC

6875): α = 11^{h}56^{m}25^{s}.87, δ = +55^{o}07′43′′.2 (equinox 2000.0). R. Arbour, S. Wonston, England, reports the following position end figures: 25^{s}.76, 40′′.1.

1998 April 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green